OpenStudy (anonymous):

I have done all the work I need someone to check Please help Iwould type the problem ASAP

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ dx }{ \sqrt{5x+8} }\] Taking U as 5x + 8 du=5 dx \[\frac{ 1 }{ 5 }\int\limits_{}^{}\frac{ 5 * dx }{ u ^{1/2} }\] \[\frac{ 1 }{ 5 }\int\limits_{}^{}u ^{-1/2} * du \] \[\frac{ 1 }{ 5 } * u ^{1/2}\]

OpenStudy (anonymous):

I don't know where i am going wrong

OpenStudy (anonymous):

@sidsiddhartha

OpenStudy (sidsiddhartha):

when ur integrating u_-1/2 then it will produce 2*u^1/2

OpenStudy (sidsiddhartha):